Johan got into the novelties business almost by accident. John often made little figurines as gifts for friends. Occasionally, he would set up a booth at flea market and sell the figurines along with jewellery. Little by little, demand for the figurines, now called jollies, grew, and he began to reproduce some of the favourite in resin, using moulds of the originals. The day came when a buyer for a major department offered him a contract to produce 1 500 figurines of various designs for R10 000,00. Johan realised that it was time to get down to business. The variable costs amounted to an average of R6,00 per unit. To produce the order, Johan would have to rent an industrial facility for a month, which would cost them R4 000,00.

a. Calculate the business operating break-even point.

b. Calculate the EBIT on the departmental store order.

c. If Johan renegotiates the contract at a price of R10,00 per figurine, what will the EBIT be?


d. If the buyer pays R8,00 per unit, but is willing to negotiate quantity, what quantity of figurines will result in an EBIT of R4 000,00?

e. At this time, lollies come in 15 different varieties. Whereas the average variable cost per unit is R6,00, the actual cost varies from unit to unit. What recommendation would you have for Johan with regard to pricing and /or the numbers and types of units that they offer for sale?
Answer #1

You have asked a question with multiple sub parts, in the same post. I have addressed the first four. Please post the balance sub parts separately.

(a) Sale Price = Revenue / Number of unit = 10,000 / 1,500 = 6.67

Let Q be the break even quantity. Hence,

(Sale Price - variable cost) x Q - Fixed cost = 0

Hence, (6.67 - 6) x Q - 4,000 = 0

Hence, break even point = Q = 4,000 / 0.67 = 6,000

(b) the EBIT on the departmental store order = Revenue - variable cost x Q - Fixed cost = 10,000 - 6 x 1,500 - 4,000 = - 3,000

(c) If Sale price is now 10 then EBIT = (Sale Price - variable cost) x Q - Fixed cost = (10 - 6) x 6,000 - 4,000 = 20,000

(d) Let N be that quantity. Hence,

(Sale Price - variable cost) x Q - Fixed cost = EBIT

Hence, (8 - 6) x Q - 4,000 = 4,000

Hence, Q = 8,000 / 2 = 4,000